OpenStudy (anonymous):

g(x)=x^2+8x+5 Does the function have a minimum or maximum value? What is the function's minimum or maximum value? Where does the minimum or maximum value occur?

OpenStudy (anonymous):

y'=2x=-8 x=-4

OpenStudy (anonymous):

I got -4 -11

OpenStudy (anonymous):

take the derivative.. g'(x) = 2x +8 then set that to zero 2x+8 = 0 x= -4 put -4 into the original equation (-4)^2 +(-4)(8) + 5 16 -32 +5 -11 has a minimum at (-4,-11)

OpenStudy (anonymous):

It's quadratic with positive coefficients, so it will have a minimum, not a maximum. The location of the minimum is at (-4,-11) as shown above.

OpenStudy (anonymous):

\[\sqrt[3]{24}\] my answer\[2\sqrt[3]{3}\]